OpenStudy (anonymous):

The formula for converting degrees Fahrenheit to degrees Celsius is: C=5/9 (F-32). Find C for F= 5 degrees

OpenStudy (alexandervonhumboldt2):

c=5/9*(5-32)=5/9*-27=-135/9=-15

OpenStudy (anonymous):

A.-27 B.5 C.-9 D.-15 They are all degrees

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

D.-15

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

Are u sure?

OpenStudy (alexandervonhumboldt2):

yeah